solve for x: -4(x-2)+4=6(x-2)-7x
x=9
solve for x: 2(x+1)-3x=3(3+2x)
x=0
solve for x: 7x=-2x-63
x=-7
solve for y: 7y-5=9y+29
y=-17
my daughter is in summer school and has 4 questions she is unsure of the answers - I am of no help - can you tell us if her answers are correct?

Answer by scott8148(6628)

-4(x-2)+4=6(x-2)-7x ___ distributing ___ -4x + 8 + 4 = 6x - 12 - 7x ___ 24 = 3x

2(x+1)-3x=3(3+2x) ___ distributing ___ 2x + 2 - 3x = 9 + 6x ___ -7 = 7x

7x=-2x-63 ___ looks good

7y-5=9y+29 ___ looks good